OpenStudy (anonymous):

integral of [e^arcsin(x)] dx /[sqrt(1-x^2)]

OpenStudy (luigi0210):

Easier than it looks ;)

OpenStudy (anonymous):

haha I can see the sqrt(1-x^2) in denominator will lead to arcsin x

OpenStudy (luigi0210):

Tell me, what's \(\Large \frac{d}{dx} ~sin^{-1}x\)

OpenStudy (anonymous):

\[1\div(\sqrt{1-x ^{2}}\]

OpenStudy (luigi0210):

Use a u-sub and get this little thing :) \[\Large \int e^u~du\] \(\Large u=arcsinx \) so that means \(\Large du=\frac{1}{\sqrt{1-x^2}} dx\)

OpenStudy (amilapsn):

you can also do this question by substituting x=sinz

OpenStudy (anonymous):

Excellent! Thank you!

OpenStudy (anonymous):

wow.. that really was easier haha

OpenStudy (luigi0210):

You're welcome~